Determine the value of the missing A/P factor for an interest rate of 19% and n=10 years, that is, (A/P,19%,10) and (b) find the value of the missing (F/P,4%,48) factor
Answer #1

a) Value of (A/P,19%,10) = [i(1 + i)n] / [(1 + i)n - 1]

                        = [0.19(1 + 0.19)10] / [(1 + 0.19)10 - 1]

                                        = 1.08199 / 4.69468

                                        = 0.2305

b) Value of (F/P, 4%, 48) = (1+ i)n

                                       = (1 + 0.04)48

                                       = 6.5705